OpenStudy (anonymous):

-5x-5y=50 Find the value of y when x=-8 .-.

OpenStudy (anonymous):

y=-2

OpenStudy (catlover5925):

@Abhisar

OpenStudy (anonymous):

-5x - 5y = 50 Plug in x. -5(-8) - 5y = 50 40 - 5y = 50 Subtract each side by 40. -5y = 10 Divide each side by -5. y = -2 <--- Answer. (: